A charge Q is divided in 2 parts. The two charges kept at a distance apart have maximum Colombian repulsion then the ratio of Q and one of the parts is given by 

Grade:12

1 Answers

Eshan
5 years ago
Dear student.

Let one of the parts bex. Then the other part isQ-x.

Hence the force between them at some distancedisF=\dfrac{x(Q-x)}{d^2}
This is maximum whenx(Q-x)=Qx-x^2is maximum.

This is maximum for\dfrac{Qx-x^2}{dx}=0
\implies Q-2x=0
\implies x=\dfrac{Q}{2}
Hence the required ratio=\dfrac{Q}{x}=2:1